Actual Fixed Costs = $300,000 Actual Variable Cost = $800,000 Budgeted Fixed Cost = $280,000 Capacity Level = 80% Variable Cost per Hour = $200 Total Hours = 10,000

Compute budget variance for fixed and variable costs
Answer #1

Variable cost variance = standard cost - Actual cost

= 10000*80%*200 - 800000 = $80000 favorable

Fixed cost variance = 280000 - 300000 = $20000 unfavorable